solid xenon forms crystals with a face centered unit cell that has an edge of 620  pm.


1. calculate the atomic radius of xenon.


2. calculate the packing efficiency if xenon.

Grade:

1 Answers

Yogita Bang
11 years ago

1) For fcc, edge = √8r

Therefore, 620 = 2.82*r

i.e., r = 620/2.82 = 219.86 pm

2) For fcc, packing efficiency = 74%